OpenStudy (fairy733189):

Which graph best represents the function f(x) = 1(1.5)x?

OpenStudy (fairy733189):

i'll attach the image

OpenStudy (bobo-i-bo):

Consider x=0

OpenStudy (fairy733189):

Hang on there's still more

OpenStudy (vishweshshrimali5):

@Fairy733189 also check whether the function is really 1(1.5)x

OpenStudy (fairy733189):

Sorry the x is an exponent

OpenStudy (vishweshshrimali5):

Great! So, \(f = 1*1.5^{x}\) ?

OpenStudy (fairy733189):

Yes!!!

OpenStudy (vishweshshrimali5):

Great! Okay so as Bobo-i-bo said, what will be value of f when x = 0 ?

OpenStudy (fairy733189):

Would it be 1 since 1.5 to the 0 power = 1 and 1 * 1 = 1

OpenStudy (vishweshshrimali5):

Perfect!

OpenStudy (vishweshshrimali5):

Now, see which of the option has y = 1 at x = 1

OpenStudy (vishweshshrimali5):

Sorry at x = 0

OpenStudy (fairy733189):

Let me check

OpenStudy (vishweshshrimali5):

Sure

OpenStudy (fairy733189):

Option 2

OpenStudy (fairy733189):

?

OpenStudy (vishweshshrimali5):

Perfect!!

OpenStudy (fairy733189):

Awesome. But am I always supposed to substitute x with 0 or can it be with any number?

OpenStudy (vishweshshrimali5):

It can be any number and you can find out the value of function at that x,... then your graph should have y = value of function you got and x = value you assumed for x...

OpenStudy (fairy733189):

Ok but I don't e=really understand it because I substituted 2 but it wasn't correct on the graph

OpenStudy (fairy733189):

Can you please explain it to me?

OpenStudy (vishweshshrimali5):

Yeah sure... we are talking of the same question? \(f = 1 * 1.5^x\)

OpenStudy (fairy733189):

yes!

OpenStudy (vishweshshrimali5):

Great! Now, when I put x = 2 => f = 1 * 1.5^2 = 1 * 2.25 = 2.25

OpenStudy (fairy733189):

Yes thats what I got but wouldn't it be like that on the graph. I just don't understand that part because when the y = 2 x would = 2.25 but the graph isn't like that

OpenStudy (vishweshshrimali5):

See at x = 2, the graph is just a bit above 2... that shows that graph is close enough.. is that actually 2.25 we can't say.. but then we are just looking for best alternative

OpenStudy (fairy733189):

Oh ok I see it now

OpenStudy (fairy733189):

Thank you so much!!!

OpenStudy (fairy733189):

I have another problem I need help with. I am going to close this one and open a new one. Can you help me with it too?

OpenStudy (vishweshshrimali5):

Yeah sure